The raft is made up of ten 18 cm diameter logs, each with a length of 200 cm. A 200cm x 200cm x 4cm wood board sits on top of the logs.

Assume the wood has a density of 600 kg/m^3

Using your sheet of paper, calculate the total mass of the raft including the logs and the wood board, in units of kg (kilogram).
Answer #1

Dimensions Of Log :- length (l) = 200 cm

Diameter (d) = 18 cm

Volume of one Cylindrical Log = (π/4)d2l = (π/4) x 182 x 200 = 50893.8 cm3 = 50893.8 x 10-6 m3

Mass of one Cylindrical Log = Density of Log Material x Volume of Log

= 600 x 50893.8 x 10-6

   = 30.54 kg

Mass of 18 Cylindrical Log = 18x30.54 kg = 549.72 kg

Dimensions Of Board :-   200 cm X 200 cm X 4 cm

Volume of Board = 200x200x4 = 160000 cm3 = 160000 x 10-6 m3

Mass of Board = Density of Board Material x Volume of Board

= 600 x 160000 x 10-6

  = 96 kg

Total Mass of the Raft Including 18 logs and Board = 549.72 + 96 = 645.72 kg

Answer :- Hence the Total Mass of the Raft Including 18 logs and Board is 645.72 kg.